OpenStudy (anonymous):

hey I'm really confused on this, i just need an answer please haha! :) Write the equation in standard form of the line through J(4, 1)with slope –4. (2 points)

OpenStudy (anonymous):

y = mx+b You know m (slope, -4) and you have a point x,y (4,1). Plug them into the equation and solve for b!

OpenStudy (jennychan12):

you have a point and the slope. point-slope form would be easier: Y-Y1 = m(X-X1) (Y1,X1) -- (4,1) so it becomes Y-1 = 4(X-4) solve for Y.